Suppose that the Sales in December 2019 were 98,674. The sales are expected to grow by 7% in January 2020, by 8% in February 2020, and by 16% in March 2020. After that, the sales are expected to grow by 5%.

The purchases are 28% of sales. The purchases will be paid in the same month. The general administrative expenses are $16,541 and will be paid in the same month. The depreciation expense is $11,926 per month. The interest expenses are $6,980 per quarter and will be paid in March, June, September, and December.

What is the cash outflow for June, 2020?

Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box.
Answer #1
Amount in dollars Dec, 19 Jan,20 Feb, 20 Mar, 20 Apr, 20 May, 20 June, 20
a) Sales 98674.00 105581.18 114027.67 132272.10 138885.71 145829.99 153121.49
b) Purchases (28% of a) 27628.72 29562.73 31927.75 37036.19 38888.00 40832.40 -42874.02
c) General Administrative expenses 16541.00 16541.00 16541.00 16541.00 16541.00 16541.00 16541.00
d) Depreciation (Non cash expense) 0.00 0.00 0.00 0.00 0.00 0.00 0.00
e) Interest Expense 6980.00 0.00 0.00 6980.00 0.00 0.00 6980.00
f) Cash outflow (b+c+d+e) 51149.72 46103.73 48468.75 60557.19 55429.00 57373.40 -19353.02
Calculation of Sales
Jan, 2019= 98674*107%=105581.18
Feb, 20=105581.18*108%=114027.67
Mar, 20=114027.67*116%=132272.10
Apr, 20=132272.10*105%=138885.71
May, 20=138885.71*105%=145829.99
Jun, 20= 145829.99*105%=153121.49